The lengths of pregnancies in a small rural village are normally distributed with a mean of 265 days and a standard deviation of 16 days.

In what range would you expect to find the middle 68% of most pregnancies?
Between  and .

If you were to draw samples of size 48 from this population, in what range would you expect to find the middle 68% of most averages for the lengths of pregnancies in the sample?
Between  and .

Enter your answers as numbers. Your answers should be accurate to 1 decimal places.
Answer #1

ogg from eph I f0 ore get K-265 16 14 265 to.ggx16 = 280-8 : The middle 68.1. of most 20g days. pregnancies age ne 48 The midwe K-265 The midolle 68% of most everages for the from eagh ® so get agg = 2-31 R = 265 t d.ggx 2.31 K= 267.3) bength of preg